# Critical behavior of the Schwinger model via gauge-invariant VUMPS

### v2

@@ II.1 Lattice-Regularized Schwinger Model
where $\sigma_{\pm}\equiv(\sigma_{x}\pm i\sigma_{y})/2$
@@
Since $L(n)\in\mathbb{Z}$, we can always restrict ourselves to the region $0\leq\theta\leq\pi$ by a suitable redefinition of $L(n)$.
@@
The lattice Hamiltonian (8) possesses some useful features. The Hamiltonian of The lattice Schwinger model defined by the Hamiltonian 8 and the Gauss law constraint (7) is invariant under translation over two sites, $T^{2}$. In addition, it is also invariant under $CT$ transformation for special values of $\theta=0$ and $\theta=\pi$, where $CT$ is the product of $T$ and the charge conjugation operator $C$ defined by
@@ IV.1 Double extrapolations
Figure 1: Plot of the simulation data for $(\delta,\epsilon_{1})$ with $ga=0.1$ and $m/g=0.333$. For $\delta<9\times 10^{-4}$ ($D>240$), $\delta$ and $\epsilon_{1}$ show a linear relationship as described in (25). The corresponding fitting line is displayed in the figure.
@@
To numerically estimate $\epsilon_{1,\infty}=\epsilon_{1,\infty}(m/g,ga)$, $(m/g)_{*}$ $=m_{*}(ga)/g$, $(m/g)_{c}$, and their uncertainties, we proceed as follows. Let us take $G\equiv\{0.1,0.2,0.3,0.4\}$. For each $ga\in G$, we choose a set $V_{ga}$ of real numbers and we run VUMPS simulations for $m/g\in V_{ga}$ for the bond dimension $D$ large enough but not exceeding 500. For each $(ga,m/g)$, we fit the numerical data representing $(\delta,\epsilon_{1})$, obtained for high enough $D$, by the linear function (25) to obtain the value of $\epsilon_{1,\infty}$. In the standard linear fit where the data $(x_{i},y_{i})$ are fitted by a linear function $y=ax+b$, the uncertainty in $b$ can be estimated from the knowledge of the uncertainty (assumed uniform, i.e., independent of $i$) in the data $y_{i}$, and the value of the reduced chi-squared should be close to 1 for a good fit. In our case, we do not know the uncertainty in the value of $\epsilon_{1}$ for given $\delta$. We apply the common procedure [78] where we choose the uncertainty in $\epsilon_{1}$ such that the reduced chi-squared is 1, and use it to estimate the uncertainty in $\epsilon_{1,\infty}$. The uncertainties are shown as error bars in FIG. 2. We then fit the data set $\{(m/g,\epsilon_{1,\infty}(m/g))\,|\,m/g\in V_{ga}\}$ by a linear function in $m/g$ in each of the two regions $m/g<(m/g)_{*}$ and $m/g>(m/g)_{*}$. [^7] Since we perform VUMPS only for a finite set of the values of $m/g$, when dividing the range of $m/g$ into the two regions, it is enough to know the value of $(m/g)_{*}$ only roughly and its $V$-dependence can be neglected. This fit gives the value of $m_{*}/g$ together with its uncertainty, for a given value of $ga$ and a choice of $V_{ga}$ (FIG. 2). Let us denote the tuple $(V_{ga=0.1},V_{0.2},V_{0.3},V_{0.4})$ by $V$. In the next step, we extrapolate to $ga=0$ by fitting $(m/g)_{*}$ with a polynomial in $ga$ and obtain the $V$-dependent estimate $m_{c,V}/g$ for the continuum critical mass and its uncertainty (FIG. 3).
@@
The above fits show that the coefficient $C_{1}$ of $ga$ is rather small, motivating us to fit the data for $(ga,m_{*}/g)$ with $C_{0}+C_{2}(ga)^{2}$. This gives the blue-star ($ga=0.1,0.2,0.3,0.4$) and brown-cross ($ga=0.1,0.2,0.3$) data points shown in FIG. 4. The procedure in the previous paragraph gives estimates for $(m/g)_{c}$ as $0.333552(1)$ and $0.333554(2)$ for the data sets, respectively. These values are smaller than but consistent with (28).
@@ IV.2 Double collapse: A finite size scaling
$\tilde{\xi}[tL^{\Delta^{\rm IR}_{t}}\Lambda^{-\Delta^{\rm UV}_{t}}]\equiv\xi/(\Lambda L),$
$\tilde{\phi}[tL^{\Delta^{\rm IR}_{t}}\Lambda^{-\Delta^{\rm UV}_{t}}]\equiv L^{\Delta^{\rm IR}_{\phi}}\Lambda^{-\Delta^{\rm UV}_{\phi}}\phi,$ (35)
@@
Figure 5: Double collapse of the correlation length (the left), the local order parameter (the middle) and the entanglement entropy (the right). The UV and IR exponents are fixed. See main text for details.
FIG. 5 shows data points of scale invariant functions eq. (35) for the corresponding optimized parameter set. One can see from each panel that all data points collapse onto a single curve. The overall behavior of double collapse in FIG. 5 is the same as that of the Euclidean $\lambda\phi^{4}$-theory [64, 65].

### v3

# Critical behavior of the Schwinger model via gauge-invariant variational uniform matrix product states

@@ II.1 Lattice-Regularized Schwinger Model
where $\sigma_{\pm}\equiv(\sigma_{x}\pm i\sigma_{y})/2$.
@@
Since $L(n)\in\mathbb{Z}$, we can always restrict ourselves to the region $0\leq\theta\leq\pi$ by a suitable redefinition of $L(n)$. We emphasize that, in contrast to many works involving MPS and quantum simulations where the electric field $L(n)$ is eliminated by solving the Gauss law constraint, our approach retains it and instead incorporates the constraint through an explicit MPS ansatz.
@@
The lattice Schwinger model defined by the Hamiltonian (8) and the Gauss law constraint (7) is invariant under translation over two sites, $T^{2}$. In addition, it is also invariant under $CT$ transformation for special values of $\theta=0$ and $\theta=\pi$, where $CT$ is the product of $T$ and the naive charge conjugation operator $C$ defined by [^4] Some works such as [70, 69] refer to our $CT$ simply as charge conjugation $C$. Here we follow [14] for $\theta=0$ and also incorporate the case $\theta=\pi$.
@@ IV.1 Double extrapolations
Figure 1: Plot of the simulation data for $(\delta,\epsilon_{1})$ with $ga=0.1$ and $m/g=0.333$. For $\delta<9\times 10^{-4}$ ($D>220$), $\delta$ and $\epsilon_{1}$ show a linear relationship as described in (25). The corresponding fitting line is displayed in the figure. For fixed $ga$ and $m/g<(m/g)_{*}$, the simulations with large $\delta$ (small $D$) prepare $CT$-breaking states rather than approximating the true $CT$-preserving ground state. See also footnote 7.
@@
To numerically estimate $\epsilon_{1,\infty}=\epsilon_{1,\infty}(m/g,ga)$, $(m/g)_{*}$ $=m_{*}(ga)/g$, $(m/g)_{c}$, and their uncertainties, we proceed as follows. Let us take $G\equiv\{0.1,0.2,0.3,0.4\}$. For each $ga\in G$, we choose a set $V_{ga}$ of real numbers and we run VUMPS simulations for $m/g\in V_{ga}$ for the bond dimension $D$ large enough but not exceeding 500. For each $(ga,m/g)$, we fit the numerical data representing $(\delta,\epsilon_{1})$, obtained for high enough $D$, by the linear function (25) to obtain the value of $\epsilon_{1,\infty}$. [^8] It is important to use $\delta(D)$ rather than $1/D$ for extrapolation. For example, the linear fit with $\delta(D)$ gives a more precise estimate of $\epsilon_{1,\infty}$ than the quadratic fit with $1/D$ by a factor of 10 for the data shown in FIG. 1. In the standard linear fit where the data $(x_{i},y_{i})$ are fitted by a linear function $y=ax+b$, the uncertainty in $b$ can be estimated from the knowledge of the uncertainty (assumed uniform, i.e., independent of $i$) in the data $y_{i}$, and the value of the reduced chi-squared should be close to 1 for a good fit. In our case, we do not know the uncertainty in the value of $\epsilon_{1}$ for given $\delta$. We apply the common procedure [79] where we choose the uncertainty in $\epsilon_{1}$ such that the reduced chi-squared is 1, and use it to estimate the uncertainty in $\epsilon_{1,\infty}$. The uncertainties are shown as error bars in FIG. 2. We then fit the data set $\{(m/g,\epsilon_{1,\infty}(m/g))\,|\,m/g\in V_{ga}\}$ by a linear function in $m/g$ in each of the two regions $m/g<(m/g)_{*}$ and $m/g>(m/g)_{*}$. [^9] Since we perform VUMPS only for a finite set of the values of $m/g$, when dividing the range of $m/g$ into the two regions, it is enough to know the value of $(m/g)_{*}$ only roughly and its $V$-dependence can be neglected. This fit gives the value of $m_{*}/g$ together with its uncertainty, for a given value of $ga$ and a choice of $V_{ga}$ (FIG. 2). Let us denote the tuple $(V_{ga=0.1},V_{0.2},V_{0.3},V_{0.4})$ by $V$. In the next step, we extrapolate to $ga=0$ by fitting $(m/g)_{*}$ with a polynomial in $ga$ and obtain the $V$-dependent estimate $m_{c,V}/g$ for the continuum critical mass and its uncertainty (FIG. 3).
@@
The above fits show that the coefficient $C_{1}$ of $ga$ is rather small, as was originally observed in [69] for a simulation with the mass shift (10). This motivates us to fit the data for $(ga,m_{*}/g)$ with $C_{0}+C_{2}(ga)^{2}$. This gives the blue-star ($ga=0.1,0.2,0.3,0.4$) and brown-cross ($ga=0.1,0.2,0.3$) data points shown in FIG. 4. The procedure in the previous paragraph gives estimates for $(m/g)_{c}$ as $0.333552(1)$ and $0.333554(2)$ for the data sets, respectively. These values are smaller than but consistent with (28).
@@ IV.2 Double collapse: A finite size scaling
$\tilde{\xi}[tL^{\Delta^{\rm IR}_{t}}\Lambda^{-\Delta^{\rm UV}_{t}}]=\xi/(\Lambda L),$
$\tilde{\phi}[tL^{\Delta^{\rm IR}_{t}}\Lambda^{-\Delta^{\rm UV}_{t}}]=L^{\Delta^{\rm IR}_{\phi}}\Lambda^{-\Delta^{\rm UV}_{\phi}}\phi,$ (35)
@@
Figure 5: Color online. Generated data points used in the double collapse on $(m/g,D_{\rm tot})$-plane (top left). Double collapse of the correlation length (top right), the local order parameter (bottom left) and the entanglement entropy (bottom right). The UV and IR exponents are fixed. See the main text for details.
In FIG. 5, we plot the three scale invariant functions in eq. (35), each with the corresponding optimized parameter set. A distinct color is assigned to each $ga$ value, and the generated data points on the $(m/g,D_{\rm tot})$-plane are also shown. One can see from each panel that all data points collapse onto a single curve. The overall behavior of double collapse in FIG. 5 is the same as that of the Euclidean $\lambda\phi^{4}$-theory [64, 65].